Operations Report – April 23rd

Crew 297 Operations Report 23-04-2024

SOL: 9

Name of person filing report: Matt Lynch (Engineer)

Non-nominal systems: Suit 2 (stuck valve), Suit 9 (broken visor), Suit 11 (fan issue),

Notes on non-nominal systems: The issues with Suit 2 and 9 were already noted before our mission. Suit 11 will no longer be used on this mission.


Opportunity rover used:

Hours: 184.2

Beginning charge: 100%

Ending charge: 50%

Currently charging: Yes

Curiosity rover used:

Hours: 279.5

Beginning charge: 100%

Ending charge: 31%

Currently charging: Yes

General notes on rovers: NA

Summary of Hab operations:

WATER USE: 29.3 gallons

Water (static tank): 204.1 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no

Summary of internet: Looking up recipes, downloading entertainment, catching up on Earthly news

Summary of suits and radios:

Radios 1,2,3,4,5, and 6 were used.

EVA Suits 1,4,5 and 6 were used.

Summary of GreenHab operations:

WATER USE: 5 gallons

Heater: Yes

Supplemental light: N/A

Harvest: 27g of parsley, 3g of cherry tomatoes

Summary of ScienceDome operations: EVA #13 had the robot operators inside the ScienceDome

Summary of RAM operations: (Tools used, work done): NA

Summary of any observatory issues: NA

Copyright © The Mars Society. All rights reserved. | Main Site